Use the given information to determine the tract trigonometric value.
sin(theta)=-1/3, 180 < theta < 270 ; tan(theta)

Answer by Alan3354(69443)

Use the given information to determine the tract trigonometric value.
sin(theta)=-1/3, 180 < theta < 270 ; tan(theta)
-----------
cos = sqrt(1-sin^2) = sqrt(8)/3
---
tan = sin/cos
tan = (1/3)/sqrt(8)/3 = 1/sqrt(8)
tan = sqrt(2)/4
----
tangent is positive in Q3